Solve: -6(x + 3)-10 = 3(-2x -5)-13
My answer (not sure if it's right):
-6x + 18 -10 = 3(-2x -5)-13
-6x + 18 -10 = -6x -15 -13
-6x + 18 -10 = -6x -15 -13
-6x + 28 = -6x -2
-6x + 28 = -6x -2
+ 6x + 6x
_________________________
x + 28 -2
- 28
__________________________
x x = -2

Answer by jim_thompson5910(35256)

-6x + 18 -10 = 3(-2x -5)-13
-6x + 18 -10 = -6x -15 -13        Distribute
-6x + 8 =-6x - 28                 Combine like terms
-6x + 8 = -6x -28                 Subtract 8 from both sides
     -8 =     -8
-6x = -6x -36                     Add 6x to both sides
+6x = +6x 

0=-36


Since this is not true, this equation has no solution
If you graphed these two equations, you will get two parallel lines.